### v1

@@ Abstract
This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing both their potential and their limitations from a computational social science perspective. The first part reviews recent findings on the ability of LLMs to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while also highlighting significant limitations such as cognitive biases, lack of true understanding, and inconsistencies in behavior. The second part surveys emerging applications of LLMs in multi-agent simulation frameworks, focusing on system architectures, scale, and validation strategies. Notable projects such as Generative Agents (Smallville) and AgentSociety are discussed in terms of their design choices, empirical grounding, and methodological innovations. Particular attention is given to the challenges of behavioral fidelity, calibration, and reproducibility in large-scale LLM-driven simulations. The final section distinguishes between contexts where LLMs, like other “black-box” systems, offer direct value—such as interactive simulations and serious games—and those where their use is more problematic, notably in explanatory or predictive modeling. The paper concludes by advocating for hybrid approaches that integrate LLMs into traditional agent-based modeling platforms (GAMA, Netlogo, etc), enabling modelers to combine the expressive flexibility of language-based reasoning with the transparency and analytical rigor of classical rule-based systems.
keywords: Large Language Models, Social Simulation, Agent-Based Modeling, Human Behavior Modeling
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a transformative class of deep learning architectures capable of generating human-like text across a wide range of domains. Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories (Vaswani 2017; Devlin 2019; Brown et al. 2020). Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation (Tonmoy 2024). Importantly, the magnitude and diversity of these training corpora mean that LLMs have effectively absorbed an immense statistical imprint of human language use across cultures, domains, and contexts. This goes far beyond learning grammar or syntax: LLMs encode traces of how people reason, argue, empathize, hesitate, and make decisions — all embedded in the linguistic fabric of their training data. Subsequently, these models are fine-tuned via instruction tuning, a supervised process in which models learn to follow natural language commands across a wide range of tasks. This is typically achieved using curated instruction datasets (Sanh et al. 2021; Wei et al. 2021), sometimes augmented with human feedback to better align model outputs with user intent (Ouyang et al. 2022).
@@
A critical area of current investigation concerns the capacity of LLMs to simulate human behavior. One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor (Jones & Bergen 2025). Recent empirical findings suggest that state-of-the-art LLMs, notably LLaMa-3.1 and GPT-4.5, have demonstrated convincing performance on a three-party version of this test. In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate (Jones & Bergen 2025). This represents a notable advancement over previous models, which had only succeeded under less rigorous two-party test conditions. These results underscore the impressive ability of modern LLMs to adopt complex linguistic styles and behavioral nuances, particularly when guided by carefully crafted persona prompts — a form of cognitive prompt engineering that remains a highly precise and non-trivial technique requiring expert design and task sensitivity.
@@
The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems (Park 2023; Anthis et al. 2025; Kirkeby-Hinrup & Stenseke 2025). However, the capacity of LLMs to produce believable human-like dialogue is not indicative of genuine understanding or consciousness. Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent (Jones & Bergen 2025). The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated (Jones & Bergen 2025), something well-known in psychology as the intentional stance (Dennett 1989). This distinction between functional imitation and authentic comprehension is particularly salient for social simulation. If LLMs merely replicate surface-level features of human behavior, their utility for generating novel insights into social processes — rather than reproducing known patterns — is fundamentally limited. Moreover, their use as proxies for human agents in scientific studies raises questions about external validity, representational fidelity, and cognitive realism. These concerns constitute a central theme of this paper and inform our critical assessment of LLMs as generative agents in artificial societies.
This position paper proceeds as follows. The first section examines the psychological plausibility of LLMs, focusing on their performance in tasks designed to evaluate theory of mind, social reasoning, and behavioral consistency. The second section surveys recent efforts to embed LLMs in agent-based social simulations, analyzing architectural strategies, scalability, and validation methodologies. The third section discusses both the potential contributions and limitations of LLM-based simulations, with particular emphasis on issues such as bias, interpretability, computational cost, and appropriate domains of application. Finally, the last section outlines future directions, including the integration of LLMs into classical modeling environments (e.g., GAMA, NetLogo) to support hybrid models that combine the explanatory clarity of rule-based approaches with the generative flexibility of LLMs.
2 LLMs in Representing Human Behavior
@@
Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as Molchanova et al. (2025), raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits. This section provides a structured review of the evidence regarding these capabilities and limitations.
@@ 2.1 Theory of Mind and Cognitive Reasoning Tasks
Theory of Mind (ToM) refers to the human capacity to attribute beliefs, intentions, desires, and knowledge to others. It is considered a foundational element of social cognition and typically emerges in early childhood. One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test (Baron-Cohen et al. 1985). In this task, a child observes a scene in which Sally places a marble in her basket and leaves the room. Anne then moves the marble to a box. Upon Sally’s return, the child is asked where she will look for her marble. A correct response—indicating Sally will look in the basket—demonstrates an understanding that Sally holds a false belief, which differs from the child’s own knowledge.
Recent studies have assessed the capacity of LLMs to pass such false-belief tasks. Kosinski (2023) tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks. While early models performed near chance levels, GPT-4 achieved over 75% accuracy, comparable to that of a 6 or 7-year-old child. Furthermore, GPT-4 demonstrated competence in second and third order ToM reasoning (e.g., “I believe that she thinks that he knows…”), with occasional success even at fifth and sixth-order levels, surpassing typical adult human performance in structured tasks.
The robustness of these findings has been explored in follow-up work. Jaiswal (2024) tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance. Performance was highly sensitive to phrasing; LLMs responded more accurately to questions using cognitive verbs (e.g., “think”) than to those phrased with action-oriented verbs (e.g., “look”), suggesting shallow linguistic pattern recognition rather than deep inference.
In addition to false-belief tasks, LLMs have been evaluated on their ability to interpret irony, indirect speech acts, and social narratives. Strachan et al. (2024) found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts. However, these results vary significantly with prompt design and context framing, raising questions about generalizability.
@@ 2.2 Emotion Representation and Affective Simulation
Emotions play a crucial role in human decision-making and social interaction. Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts (Lee et al. 2024). For example, given a prompt describing a tragic event, GPT-4 can produce empathetic responses and plausible emotional narratives. Nonetheless, this capacity is fundamentally linguistic. LLMs do not experience emotions nor maintain internal emotional states. Their affective responses are generated through statistical associations learned from text corpora containing human emotional expression.
@@
Studies have attempted to evaluate LLMs’ performance on emotional recognition tasks. Lai et al. (2025) report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts. Yet these capabilities depend on typical lexical cues and falter when faced with indirect or culturally nuanced expressions. Emotional modeling in LLMs remains limited to surface-level mimicry, lacking integration with motivational systems or physiological grounding, both of which are central to human affective behavior.
2.3 Bias, Representational Gaps, and Cognitive Illusions
@@
Because LLMs are trained on large-scale human-generated data, they inevitably inherit societal biases embedded in those texts. These include gender, racial, and cultural stereotypes. Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior(Santurkar 2025). In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations (Kotek et al. 2024).
@@
Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans.(Kim et al. 2025) found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks. Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—did not consistently reduce these biases. In some instances, reasoning chains amplified them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them.
Efforts to debias LLMs using prompt engineering or few-shot learning strategies have shown promise but remain partial solutions. Schramowski (2024) argues that mitigating bias in LLMs requires not only refined prompting but also upstream control over training data composition and filtering mechanisms.
2.4 Limitations in Consistency, Hallucination, and Generalization
A pervasive challenge in LLM behavior is the issue of hallucination, where the model generates outputs that are linguistically fluent but factually incorrect or inconsistent with the input (Tonmoy 2024). This phenomenon undermines the credibility of LLM-generated reasoning and introduces substantial risks in simulation contexts. Combined with susceptibility to input phrasing and lack of state consistency over time, hallucination further limits the reliability of LLMs for modeling human behavior dynamically.

### v2

@@ Abstract
This position paper examines the use of Large Language Models (LLMs) in social simulation, analyzing their potential and limitations from a computational social science perspective. We first review recent findings on LLMs’ ability to replicate key aspects of human cognition, including Theory of Mind reasoning and social inference, while identifying persistent limitations such as cognitive biases, lack of grounded understanding, and behavioral inconsistencies. We then survey emerging applications of LLMs in multi-agent simulation frameworks, examining system architectures, scalability, and validation strategies. Projects such as Generative Agents (Smallville) and AgentSociety are analyzed with respect to their empirical grounding and methodological design. Particular attention is given to the challenges of behavioral fidelity, calibration, and reproducibility in large-scale LLM-driven simulations. Finally, we distinguish between contexts where LLM-based agents provide operational value—such as interactive simulations and serious games—and contexts where their use raises epistemic concerns, particularly in explanatory or predictive modeling. We argue that hybrid approaches integrating LLMs into established agent-based modeling platforms such as GAMA and NetLogo may offer a promising compromise between expressive flexibility and analytical transparency. Building on this analysis, we outline a conceptual research direction termed Hybrid Constitutional Architectures, which proposes a stratified integration of classical agent-based models (ABMs), small language models (SLMs), and LLMs within established platforms such as GAMA and NetLogo.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a transformative class of deep learning architectures capable of generating human-like text across a wide range of domains. Built upon the transformer architecture, which employs parallelized attention mechanisms rather than sequential processing, these models are pre-trained on massive corpora comprising web data, encyclopedic content, and code repositories [74, 20, 11]. Through unsupervised or self-supervised learning, LLMs acquire syntactic, semantic, and pragmatic language competencies, enabling them to perform a broad array of downstream tasks such as question answering, summarization, translation, and code generation [73]. Importantly, the magnitude and diversity of these training corpora mean that LLMs have effectively absorbed an immense statistical imprint of human language use across cultures, domains, and contexts. This goes far beyond learning grammar or syntax: LLMs encode traces of how people reason, argue, empathize, hesitate, and make decisions — all embedded in the linguistic fabric of their training data. Subsequently, these models are fine-tuned via instruction tuning, a supervised process in which models learn to follow natural language commands across a wide range of tasks. This is typically achieved using curated instruction datasets [58, 79], sometimes augmented with human feedback to better align model outputs with user intent [53].
Given this capability to emulate human-like reasoning patterns, LLMs have naturally attracted significant attention within the Agent-Based Modeling (ABM) community. However, as their integration into this field accelerates, it is crucial to distinguish between two emerging uses of Large Language Models in computational social science. The first involves using LLMs as coding assistants to facilitate the technical implementation of simulation models (e.g., generating NetLogo or GAMA code). The second, which is the sole focus of this paper, is the use of LLMs as the cognitive architecture of the agents themselves—often termed ’Generative Agents.’ We examine how LLMs can replace or augment traditional rule-based cognition to simulate human-like decision-making, social interaction, and reasoning, rather than how they assist in the software engineering process.
@@
Focusing specifically on this role as a cognitive engine, a critical area of current investigation concerns the capacity of LLMs to authentically simulate human behavior. One classical benchmark in this domain is the Turing Test, proposed by Alan Turing in 1950 as the imitation game, which evaluates whether a machine can produce conversational behavior indistinguishable from that of a human interlocutor [35]. Recent empirical findings suggest that state-of-the-art LLMs, notably LLaMa-3.1 and GPT-4.5, have demonstrated convincing performance on a three-party version of this test. In these evaluations, LLaMa-3.1 was perceived as human in 56% of cases, while GPT-4.5 achieved a 73% success rate [35]. This represents a notable advancement over previous models, which had only succeeded under less rigorous two-party test conditions. These results underscore the impressive ability of modern LLMs to adopt complex linguistic styles and behavioral nuances, particularly when guided by carefully crafted persona prompts — a form of cognitive prompt engineering that remains a highly precise and non-trivial technique requiring expert design and task sensitivity.
@@
The apparent success of LLMs in the Turing Test has naturally prompted speculation about their suitability for simulating human agents within computational social systems [54, 4, 39]. However, the capacity of LLMs to produce believable human-like dialogue is not indicative of genuine understanding or consciousness. Rather, their performance reflects advanced statistical pattern recognition — the ability to mimic linguistic structures convincingly — without necessarily modeling the underlying cognitive or emotional states of a human agent [35]. The tendency of human evaluators to anthropomorphize fluent language exacerbates this illusion, as observers may attribute intentionality or mental states to outputs that are, in essence, probabilistically generated [35], something well-known in psychology as the intentional stance [19]. This distinction between functional imitation and authentic comprehension is particularly salient for social simulation. If LLMs merely replicate surface-level features of human behavior, their utility for generating novel insights into social processes — rather than reproducing known patterns — is fundamentally limited. Moreover, their use as proxies for human agents in scientific studies raises questions about external validity, representational fidelity, and cognitive realism. These concerns constitute a central theme of this paper and inform our critical assessment of LLMs as generative agents in artificial societies.
This position paper proceeds as follows. Section 2 examines the psychological plausibility of LLMs as individual agents, evaluating their performance in social reasoning tasks while highlighting intrinsic cognitive limitations—such as hallucinations, inconsistency, and statistical biases—that challenge their reliability. Section 3 surveys the state of the art in LLM-based social simulation, reviewing emerging platforms, cognitive architectures for multi-agent systems, and current methodological strategies for empirical validation. Section 4 critically discusses the methodological barriers to scaling these agents into complex simulations, with particular emphasis on the ”micro-macro validity gap” the loss of epistemic control due to model opacity, and the ethical risks of automation bias. Finally, Section 5 proposes a synthesis of future research directions, examining the trade-offs between data-driven behavioral fine-tuning and theory-driven cognitive alignment. It advocates for ’Hybrid Constitutional Architectures’— a stratified paradigm that combines Small Language Models (SLMs) with theoretical psychological grounding—and discusses their integration into classical modeling environments (e.g., GAMA, NetLogo) to ensure groundedness and causal validity, particularly in data-scarce and culturally diverse contexts.
To ensure a comprehensive coverage of this rapidly evolving field, the literature selection for this position paper combines a systematic search with targeted exploratory retrieval. First, a structured query was conducted on the Scopus database (cutoff date: February 12) using the following boolean string: TITLE-ABS-KEY ((”Large Language Model*” OR ”Generative Agent*”) AND (”Social Simulation” OR ”Agent-Based”)). This initial search yielded a corpus of recent articles (2023-2026), from which we selected seminal works based on citation count and relevance to social science applications. Second, given the interdisciplinary and fast-paced nature of the domain (where many breakthroughs appear as preprints), we complemented this systematic extraction with LLM-assisted retrieval (using Gemini 3 Pro). This model were used to identify specific niche applications and to uncover cross-disciplinary links (e.g., cognitive psychology references) that keyword-based searches might miss. Crucially, human verification remained the final filter: every reference suggested by AI tools was manually reviewed by the authors to ensure scientific rigor, correct attribution, and contextual relevance to the arguments presented.
2 LLMs in Simulating Human Behavior: Capabilities and Cognitive Limitations
@@
Thus, while LLMs can convincingly mimic language use and pass superficial behavioral tests, the extent to which their outputs align with actual human cognition, reasoning patterns, emotional expression, and social bias remains contested.For instance, recent studies, such as [50], raise important concerns about the depth and consistency of LLMs’ psychological realism by exploring their potential to simulate stable and diverse personality traits. This section provides a structured review of the evidence regarding these capabilities and limitations.
@@ 2.1 Theory of Mind and Cognitive Reasoning Tasks
Theory of Mind (ToM) refers to the human capacity to attribute beliefs, intentions, desires, and knowledge to others. It is considered a foundational element of social cognition and typically emerges in early childhood. One of the canonical tests to assess ToM in developmental psychology is the Sally-Anne test [7]. In this task, a child observes a scene in which Sally places a marble in her basket and leaves the room. Anne then moves the marble to a box. Upon Sally’s return, the child is asked where she will look for her marble. A correct response—indicating Sally will look in the basket—demonstrates an understanding that Sally holds a false belief, which differs from the child’s own knowledge.
Recent studies have assessed the capacity of LLMs to pass such false-belief tasks. [40] tested multiple LLMs, including GPT-3.5 and GPT-4, on 40 ToM problems designed to mirror psychological benchmarks. While early models performed near chance levels, GPT-4 achieved over 75% accuracy, comparable to that of a 6 or 7-year-old child. Furthermore, GPT-4 demonstrated competence in second and third order ToM reasoning (e.g., “I believe that she thinks that he knows…”), with occasional success even at fifth and sixth-order levels, surpassing typical adult human performance in structured tasks.
The robustness of these findings has been explored in follow-up work. [32] tested LLMs on modified ToM tasks involving less prototypical narratives—such as a cat hiding a key or a user renaming folders offline—and found a significant drop in model performance. Performance was highly sensitive to phrasing; LLMs responded more accurately to questions using cognitive verbs (e.g., “think”) than to those phrased with action-oriented verbs (e.g., “look”), suggesting shallow linguistic pattern recognition rather than deep inference.
In addition to false-belief tasks, LLMs have been evaluated on their ability to interpret irony, indirect speech acts, and social narratives. [67] found that GPT-4 matched or exceeded average human performance in recognizing indirect requests, detecting sarcasm, and attributing communicative intent in narrative contexts. However, these results vary significantly with prompt design and context framing, raising questions about generalizability.
@@ 2.2 Emotion Representation and Affective Simulation
Emotions play a crucial role in human decision-making and social interaction. Unlike rule-based agents or earlier AI systems, LLMs can generate emotionally appropriate language in many contexts [45]. For example, given a prompt describing a tragic event, GPT-4 can produce empathetic responses and plausible emotional narratives. Nonetheless, this capacity is fundamentally linguistic. LLMs do not experience emotions nor maintain internal emotional states. Their affective responses are generated through statistical associations learned from text corpora containing human emotional expression.
@@
Studies have attempted to evaluate LLMs’ performance on emotional recognition tasks. [42] report that LLMs can often classify emotions expressed in text (e.g., anger, sadness) and match tones to situational prompts. Yet these capabilities depend on typical lexical cues and falter when faced with indirect or culturally nuanced expressions. Emotional modeling in LLMs remains limited to surface-level mimicry, lacking integration with motivational systems or physiological grounding, both of which are central to human affective behavior.
2.3 Anthropomorphic and Statistical Biases
2.3.1 Societal and Representational Biases
@@
Because LLMs are trained on large-scale human-generated data, they inevitably inherit and can perpetuate societal biases embedded in those texts. These include explicit and implicit stereotypes related to race, gender, religion, and culture. Studies have shown that LLMs reproduce associations consistent with biased representations: for instance, associating certain professions with male pronouns or producing racially skewed descriptions of criminal behavior [60]. In addition, structural imbalances in data availability mean that voices from underrepresented linguistic or social groups are poorly modeled, resulting in inequitable performance across different populations [41]. This ”black-box” inheritance of bias persists despite instruction tuning, and differences between model versions can lead to inconsistent user experiences and hidden harms in downstream applications [24].
2.3.2 Cognitive Biases and the Reasoning Paradox
@@
Beyond representational biases, LLMs also exhibit patterns consistent with cognitive biases observed in humans—defined as systematic deviations from rationality in judgment. [38] found that LLMs replicate anchoring effects, status quo bias, framing effects, and recency bias when posed with structured reasoning tasks. While debiasing efforts using prompt engineering have shown promise, they remain partial solutions [62]. Alarmingly, the tendency to apply ”chain-of-thought” prompting—a method designed to enhance logical reasoning—does not consistently prevent these biases. In some instances, reasoning chains actually amplify them, indicating that LLM ”reasoning” may reinforce pre-trained associations rather than correct them [38]. [62] argues that mitigating these issues requires upstream control over training data composition rather than just refined prompting.
2.3.3 Convergence towards the ”Average Persona”
Another bias that raises significant ethical issues is the convergence towards an “average persona.” LLMs fundamentally operate on statistical probabilities, predicting the ”next likely word” based on patterns in their training data rather than deep cognitive understanding [75, 57]. Consequently, while human decisions are shaped by a rich tapestry of lived experiences and future goals, an LLM merely calculates the most mathematically expected response. This results in a flattening of diversity that can inadvertently erase the characteristics of minority subgroups in simulated populations [81, 24]. Empirical evidence highlights this homogeneity in economic games, where LLMs often produce highly uniform responses compared to the wider range of choices observed in humans [81]. This trend extends to the political sphere, where models tend to produce narrower distributions of opinions that overrepresent the values of wealthy, young, and politically liberal individuals—the so-called ”WEIRD” demographic bias [59, 29].
2.3.4 Artifactual Biases from Safety Alignment
Beyond this statistical convergence, the deliberate ”safety alignment” of LLMs—achieved through Reinforcement Learning from Human Feedback (RLHF)—introduces a further critical risk: the generation of ”false negatives” in exploratory research. While training models to be ”helpful and harmless” is essential for commercial products, this normative alignment acts as a systematic filter that ”sanitizes” agent behaviors. In the context of social simulation, an over-aligned model may exhibit a ”sycophancy bias” [63], refusing to manifest conflictual, aggressive, or ”dark” social dynamics even when such behaviors are realistic and essential for understanding phenomena like polarization or radicalization. Consequently, researchers risk exploring a truncated solution space where catastrophic but plausible social outcomes are suppressed, leading to a dangerous underestimation of systemic risks. While context-rich prompts can mitigate this convergence, this variability remains contingent on the user’s intervention rather than an inherent capacity of the model for autonomous diversity [5].
2.3.5 Bounded Rationality vs. Artifactual Bias
Finally, a critical theoretical distinction must be made regarding the very definition of ’bias’ in generative social science. Following Herbert Simon’s theory of Bounded Rationality [65], human decision-making is inherently constrained by cognitive limitations and heuristics. In social simulation, these human deviations from pure logic are not defects to be removed, but constitutive features necessary for realism. Therefore, we must rigorously distinguish between ’Desirable Cognitive Biases’—such as loss aversion or emotional reactivity, which reflect authentic human irrationality—and ’Artifactual Model Biases’ stemming from the LLM’s training process. The latter include phenomena like ’Sycophancy’ (the tendency to agree with the user to be ’helpful’) [63] or ’Safety Sanitization’ (the refusal to simulate conflictual interactions). While Simon’s bounded rationality is a target to be reached, these artifactual biases are ’machine-like’ distortions that flatten social dynamics and must be actively mitigated through architectural constraints.
2.4 Intrinsic Instability: Hallucinations and Inconsistency
A pervasive challenge in LLM behavior is the issue of hallucinations, where the model generates outputs that are baseless, untruthful, or diverge from factual reality, even while maintaining a coherent linguistic structure [73]. These errors often stem from flawed training data or specific architectural biases [48, 31, 28]. Crucially, when LLMs hallucinate, they typically maintain an ’aura of authority’, generating responses that appear coherent and well-formed. This deceptive coherence creates a ’validity illusion’ that can mislead users—especially those lacking domain expertise—and introduces substantial risks in scientific applications where factual accuracy is paramount [64, 68].